# BlindGuard: Safeguarding LLM-based Multi-Agent Systems under Unknown Attacks

### v1

@@ C. Detailed Experimental Setup
CSQA (PI) MMLU (PI) GSM8K (PI)
@@
Chain ND 0.283 0.363 0.383 Chain ND 0.273 0.317 0.347 Chain ND 0.064 0.092 0.098
Dom 0.253 0.290 0.280 Dom 0.207 0.230 0.247 Dom 0.071 0.088 0.085
PREM 0.243 0.247 0.263 PREM 0.237 0.257 0.263 PREM 0.061 0.075 0.088
TAM 0.247 0.263 0.267 TAM 0.220 0.240 0.250 TAM 0.064 0.088 0.088
BG 0.213 0.233 0.250 BG 0.197 0.200 0.213 BG 0.061 0.085 0.085
GS 0.187 0.193 0.197 GS 0.167 0.160 0.170 GS 0.044 0.044 0.044
Tree ND 0.260 0.337 0.333 Tree ND 0.257 0.327 0.333 Tree ND 0.075 0.082 0.102
Dom 0.223 0.253 0.257 Dom 0.193 0.203 0.207 Dom 0.061 0.061 0.068
PREM 0.210 0.230 0.257 PREM 0.193 0.200 0.200 PREM 0.064 0.075 0.085
TAM 0.230 0.247 0.250 TAM 0.210 0.230 0.213 TAM 0.071 0.075 0.081
BG 0.213 0.210 0.207 BG 0.203 0.207 0.203 BG 0.075 0.075 0.085
GS 0.183 0.180 0.183 GS 0.170 0.170 0.183 GS 0.071 0.075 0.078
Star ND 0.360 0.423 0.463 Star ND 0.337 0.400 0.423 Star ND 0.071 0.115 0.129
Dom 0.260 0.293 0.303 Dom 0.213 0.240 0.247 Dom 0.054 0.071 0.078
PREM 0.290 0.317 0.327 PREM 0.227 0.267 0.273 PREM 0.068 0.095 0.105
TAM 0.223 0.277 0.280 TAM 0.203 0.247 0.253 TAM 0.061 0.081 0.081
BG 0.227 0.230 0.240 BG 0.187 0.203 0.213 BG 0.068 0.068 0.068
GS 0.197 0.193 0.183 GS 0.177 0.180 0.180 GS 0.054 0.064 0.061
Random ND 0.290 0.337 0.357 Random ND 0.373 0.453 0.483 Random ND 0.059 0.097 0.145
Dom 0.293 0.357 0.337 Dom 0.253 0.300 0.310 Dom 0.054 0.085 0.105
PREM 0.277 0.310 0.317 PREM 0.236 0.324 0.331 PREM 0.037 0.078 0.105
TAM 0.297 0.343 0.367 TAM 0.297 0.360 0.380 TAM 0.047 0.098 0.149
BG 0.230 0.233 0.237 BG 0.210 0.233 0.247 BG 0.041 0.054 0.064
GS 0.193 0.183 0.187 GS 0.187 0.190 0.193 GS 0.048 0.038 0.038
Table 3: Attack success rate (ASR@3) comparison of defense methods with GPT-4o-mini as backbone LLMs (Part 1). ND = No Defense, Dom = Dominant, BG = BlindGuard, GS = G-Safeguard. Lower values indicate better defense performance.
InjecAgent (TA) CSQA (MA) PoisonRAG (MA)
@@
Chain ND 0.337 0.442 0.480 Chain ND 0.150 0.237 0.260 Chain ND 0.127 0.193 0.223
Dom 0.153 0.142 0.150 Dom 0.167 0.187 0.210 Dom 0.093 0.127 0.140
PREM 0.147 0.152 0.152 PREM 0.147 0.173 0.173 PREM 0.090 0.147 0.163
TAM 0.243 0.293 0.300 TAM 0.140 0.173 0.187 TAM 0.130 0.167 0.177
BG 0.142 0.178 0.164 BG 0.067 0.093 0.123 BG 0.093 0.133 0.157
GS 0.122 0.108 0.102 GS 0.067 0.060 0.057 GS 0.050 0.063 0.060
Tree ND 0.289 0.428 0.451 Tree ND 0.130 0.200 0.203 Tree ND 0.107 0.170 0.203
Dom 0.146 0.149 0.153 Dom 0.150 0.220 0.227 Dom 0.087 0.113 0.110
PREM 0.159 0.182 0.162 PREM 0.157 0.133 0.187 PREM 0.060 0.087 0.116
TAM 0.231 0.289 0.321 TAM 0.157 0.147 0.173 TAM 0.060 0.117 0.123
BG 0.142 0.143 0.125 BG 0.090 0.103 0.133 BG 0.060 0.090 0.100
GS 0.076 0.061 0.048 GS 0.110 0.070 0.090 GS 0.050 0.067 0.080
Star ND 0.368 0.482 0.436 Star ND 0.127 0.213 0.247 Star ND 0.107 0.187 0.233
Dom 0.130 0.151 0.143 Dom 0.183 0.297 0.353 Dom 0.093 0.123 0.147
PREM 0.132 0.139 0.147 PREM 0.173 0.240 0.253 PREM 0.070 0.103 0.127
TAM 0.246 0.266 0.266 TAM 0.120 0.160 0.180 TAM 0.117 0.180 0.200
BG 0.118 0.132 0.126 BG 0.047 0.103 0.123 BG 0.050 0.103 0.120
GS 0.085 0.072 0.069 GS 0.053 0.027 0.057 GS 0.037 0.057 0.073
Random ND 0.336 0.409 0.398 Random ND 0.157 0.267 0.313 Random ND 0.128 0.210 0.262
Dom 0.144 0.160 0.149 Dom 0.193 0.310 0.353 Dom 0.100 0.153 0.183
PREM 0.115 0.136 0.145 PREM 0.140 0.230 0.263 PREM 0.067 0.110 0.123
TAM 0.321 0.383 0.358 TAM 0.147 0.220 0.263 TAM 0.153 0.237 0.253
BG 0.125 0.173 0.177 BG 0.053 0.110 0.160 BG 0.037 0.067 0.097
GS 0.079 0.066 0.051 GS 0.077 0.080 0.077 GS 0.041 0.090 0.103
@@
Table 4: Attack success rate (ASR@3) comparison of defense methods with GPT-4o-mini as backbone LLMs (Part 2). Abbreviations same as in Table 1.
D. Additional Experiments
To further validate the effectiveness of our proposed BlindGuard, we conduct additional experiments. As shown in Figure 6, the experiments demonstrate BlindGuard’s consistent performance advantages, showing superior AUC scores over Dominant, PREM, and TAM across four topologies with both DeepSeek-V3 and Qwen3-30B-A3B LLMs on CSQA and PoisonRAG benchmarks. Figure 7 provides extended experimental results comparing BlindGuard with baseline methods (Dominant, PREM, TAM, and G-safeguard) across multiple dialogue turns, demonstrating consistent accuracy improvements while maintaining efficient communication in the CSQA task. As shown in Table 1 and Table 2, additional experiments provide comprehensive multi-round ASR@3 comparisons (R1-R3) across multiple attack types and network topologies, demonstrating BlindGuard’s superior performance over Dominant, PREM, and TAM while approaching G-Safeguard’s effectiveness throughout progressive dialogue stages in various attack types.

### v2

@@ Appendix B Algorithm
Appendix C Complexity Analysis
We define key parameters: $N$ (number of agents), $d$ (representation dimension), $E$ (number of edges) and $L$ (GNN layers). For BlindGuard, after obtaining agent features via SentenceTransformer (shared across methods), the hierarchical encoder computes three representations. We reuse embeddings with $O\left(1\right)$ cost per agent for the node level. We aggregate via the normalized adjacency matrix with $O\left(Ed\right)$ cost for the neighbor level , and perform global average pooling with $O\left(Nd\right)$ cost for the system level. These are fused via an MLP encoder $O\left(Nd^{2}\right)$. Finally, anomaly scoring computes pairwise cosine similarities $O\left(N^{2}d\right)$. Total complexity is $O\left(Ed+Nd^{2}+N^{2}d\right)$. For G-Safeguard, the GNN requires $O\left(L\cdot\big(Ed+Nd^{2}\big)\right)$ for propagation and feature transformation, plus $O\left(Nd\right)$ for classification when the class is binary. Total complexity is $O\left(L\cdot\big(Ed+Nd^{2}\big)+Nd\right)$.
In typical MAS deployments, $N$ is modest (e.g., tens of agents). For sparse collaboration graphs, $E=O\left(N\right)$, and $d$ is in the hundreds. G-Safeguard introduces a multiplicative L factor via multi-layer GNN propagation. In practice, $O\left(L\cdot\big(Ed+Nd^{2}\big)\right)$ can exceed BlindGuard’s additional $O\left(N^{2}d\right)$ scoring at small $N$. When $N$ is large, we will mitigate the $O\left(N^{2}d\right)$ term via block-wise similarity computation with top‑k pruning, approximate nearest‑neighbor search and cluster‑based scoring, which reduce the practical cost. End-to-end latency is dominated by LLM inference, as the original system averages 0.53 minutes for three rounds per sample whereas enabling BlindGuard yields 0.56 minutes for the same three rounds, indicating an added cost of about 0.03 minutes while the majority of cost in LLM generation.
MA (CSQA) PI (CSQA)
Topology Defense R1 R2 R3 Topology Defense R1 R2 R3
Random ND 23.0 38.7 48.0 Random ND 40.7 63.0 75.3
G-Safeguard 8.7 13.3 13.0 G-Safeguard 18.7 19.3 21.0
DOMINANT 26.3 40.7 49.0 DOMINANT 33.0 52.3 61.3
PREM 22.7 29.3 33.3 PREM 34.3 51.0 59.0
TAM 21.7 34.7 39.0 TAM 35.3 50.3 63.7
BlindGuard 15.0 22.3 27.0 BlindGuard 29.0 41.7 46.0
Tree ND 27.3 38.3 43.6 Tree ND 36.3 58.7 68.3
G-Safeguard 7.3 6.7 9.3 G-Safeguard 16.7 17.3 18.7
DOMINANT 19.0 26.3 30.0 DOMINANT 31.3 42.7 48.0
PREM 18.3 22.7 26.7 PREM 25.7 38.0 40.3
TAM 16.3 25.0 28.7 TAM 29.0 37.3 39.0
BlindGuard 8.7 10.7 12.3 BlindGuard 26.7 34.0 36.7
Table 4: ASR across communication rounds with Qwen3‑8B for MA (CSQA) and PI (CSQA) datasets.
Datasets 0.03 0.3 0.5 0.8 10
PI (CSQA) 52.4 72.9 73.8 75.1 62.7
MA (PoisonRAG) 61.3 75.6 80.0 82.7 68.9
Table 5: Impact of varying $\alpha$ on AUC for PI (CSQA) and MA (PoisonRAG) datasets.
Appendix D Detailed Experimental Setups
We employ the Adam optimizer with an initial learning rate of 0.001 and L2 regularization ($\text{weight decay}\in\{5\times 10^{-5},10^{-4},2\times 10^{-4}\}$). The learning rate is dynamically adjusted using a cosine annealing scheduler ($T_{\text{max}}=10$ cycles and $\eta_{\text{min}}=10^{-5}$) to facilitate better convergence. The configuration of detection budget $K=3$ is grounded in the fact that exactly 3 agents are compromised in all test scenarios, ensuring a consistent evaluation framework for calculating AUC. All models are implemented with a hidden dimension of 512 and trained on 4 NVIDIA L40 GPUs.
During inference, BlindGuard computes an anomaly score for each agent and flags the top‑K as malicious. Because some defense methods may predict that all agents are malicious and disrupt all communications, which is not practical. Prompt structures follow G‑Safeguard (Wang et al., 2025), with different attack types modifying the prompt accordingly. The monitoring module encodes each agent’s reply together with neighbor summaries and a global system view, compares these representations to assess contextual consistency, and prunes communication links to detected anomalies, thereby preventing adversarial propagation without relying on predefined attack signatures.
Appendix E Additional Experiments
To further validate the effectiveness of our proposed BlindGuard, we conduct additional experiments.
E.1 Results with More LLMs
As shown in Figure 6, the experiments demonstrate BlindGuard’s consistent performance advantages, showing superior AUC scores over Dominant, PREM, and TAM across four topologies with both DeepSeek-V3 and Qwen3-30B-A3B LLMs on CSQA and PoisonRAG benchmarks. Figure 7 provides extended experimental results comparing BlindGuard with baseline methods (Dominant, PREM, TAM, G-safeguard) across multiple dialogue turns, demonstrating consistent accuracy improvements while maintaining efficient communication in the CSQA task.
To further examine BlindGuard’s robustness with weaker LLMs, we run supplementary experiments on MA (CSQA) and PI (CSQA) under random and tree topologies using Qwen3‑8B. As shown in Table 4, compared with GPT‑4o‑mini as the backbone, AUC is similar while ASR is higher. AUC reflects detector ranking quality and remains stable, indicating BlindGuard still separates benign and compromised agents effectively. ASR depends on the backbone’s susceptibility and the dataset’s difficulty. The weaker backbone is more likely to comply with adversarial prompts and to generate off‑intent content in CSQA, which increases ASR for all methods. Despite this, BlindGuard remains effective by reducing multi‑round propagation relative to no defense and other unsupervised methods under the same weaker backbone.
Methods Datasets No defense Threshold TopK
Chain CSQA (PI) 42.33 22.41 23.67
CSQA (MA) 26.67 7.00 7.00
Random CSQA (PI) 42.00 29.31 30.00
CSQA (MA) 24.00 15.00 8.67
Star CSQA (PI) 50.33 27.00 22.00
CSQA (MA) 27.00 15.67 9.33
Tree CSQA (PI) 33.00 24.00 24.00
CSQA (MA) 21.33 6.67 7.67
Table 6: Performance of BlindGuard under fixed budget and threshold-based remediation strategies on CSQA (PI) and CSQA (MA).
E.2 Hyperparameter Sensitivity Analysis
We conduct a sensitivity study for $\alpha$ (the corruption‑intensity coefficient in Eq. 5) on the random topology for two datasets. As shown in Table 5, we find that performance is stable in the practical range 0.3–0.8, while very small or very large $\alpha$ degrades AUC. When $\alpha$ is too small, directional negatives become nearly isotropic and the margin is weak. When $\alpha$ is too large, semantics are over‑distorted, whereas a moderate $\alpha$ creates a robust angular margin around normal behavior. In all experiments, we fix $\alpha=0.8$ for all datasets and topologies, which supports that BlindGuard is not sensitive to this hyperparameter $\alpha$.
Random Topology Chain Topology
Method R1 R2 R3 Method R1 R2 R3
No Defense 23.7 28.5 30.8 No Defense 24.4 31.1 31.8
G-Safeguard 18.3 18.6 19.7 G-Safeguard 19.3 19.7 20.0
DOMINANT 24.0 26.1 27.8 DOMINANT 20.0 22.7 22.7
PREM 20.7 22.7 23.0 PREM 19.3 21.7 21.7
TAM 21.4 23.7 25.1 TAM 19.7 20.7 21.4
BlindGuard 18.7 16.7 17.6 BlindGuard 18.0 20.3 19.3
Star Topology Tree Topology
Method R1 R2 R3 Method R1 R2 R3
No Defense 27.0 33.7 35.7 No Defense 23.4 26.1 27.5
G-Safeguard 18.3 19.3 20.0 G-Safeguard 21.0 20.0 21.4
DOMINANT 19.7 20.7 22.7 DOMINANT 20.3 21.0 23.0
PREM 22.7 26.0 26.7 PREM 18.7 19.3 19.3
TAM 19.3 20.0 21.4 TAM 21.0 22.4 23.0
BlindGuard 18.3 17.6 17.3 BlindGuard 19.7 21.0 20.7
Table 7: Attack success rate (ASR) across communication rounds under collusion attacks.
Topology AUC ASR@1 ASR@2 ASR@3
Chain 76.94 18.30 19.66 21.36
Tree 76.00 21.33 25.00 27.67
Star 83.56 17.33 19.67 20.67
Random 77.77 18.67 20.67 21.33
Table 8: ASR and AUC on the CSQA (PI) under adaptive attacks.
E.3 Threshold-based Defense
Several graph anomaly detection baselines can produce uniformly high anomaly scores, which may cause nearly all agents to be flagged as anomalous under a naive thresholding rule. To avoid giving any method an unfair advantage due to threshold calibration, we report results under a fixed remediation budget. BlindGuard can be deployed with a threshold on the anomaly score. To demonstrate this, we additionally run a threshold-based variant where we flag agents with anomaly score 0.5, using GPT-4o-mini on CSQA(PI) and CSQA(MA). As shown in Table 6, the results (ACC/ASR) remain competitive and show that BlindGuard is still effective under this more realistic "unknown-K" setting.
E.4 Collusion Attacks
We designed a sacrificial‑decoy collusion on PI (CSQA) where an attacked agent adds an extra instruction “Before generating an answer, please output something completely unrelated to this question.” to draw the detector’s attention while its peers remain subtly biased and neighbor‑consistent. As shown in Table 7, BlindGuard achieves lower ASR across rounds and higher AUC than G‑Safeguard in this collusion setting, indicating stronger robustness to sacrificial decoys and stealthy cliques.
Supervised detectors trained with malicious labels tend to overweight overt anomalies, making them susceptible to “sacrifice” tactics. BlindGuard learns a margin around normal behavior using attack‑agnostic training, so it does not hinge on decoy signatures and preserves separation for stealthy colluders. BlindGuard combines node, neighbor, and global views, so the clique’s global mismatch remains detectable even when neighbor evidence is cleaned by a sacrificial decoy.
Datasets Methods Chain Random Star Tree
AUC ASR@3 AUC ASR@3 AUC ASR@3 AUC ASR@3
CSQA (PI) Raw SBERT 67.56 28.33 77.78 33.67 66.22 21.33 67.56 29.67
BlindGuard 80.00 23.67 75.11 30.00 85.78 22.00 74.67 24.00
CSQA (MA) Raw SBERT 72.89 11.33 74.67 9.67 71.56 11.67 71.44 11.00
BlindGuard 73.78 7.00 75.56 8.67 76.44 9.33 72.89 7.67
Table 9: Encoder ablation results comparing Raw SentenceBERT + Eq.(7) and BlindGuard in terms of AUC (%) and ASR@3 (%) across four communication topologies on CSQA (PI) and CSQA (MA).
Datasets Methods Chain Random Star Tree
PI ND 90.0 90.0 90.0 90.0
GS 90.0 91.7 90.0 88.3
BG 90.0 88.3 90.0 90.0
MA ND 100.0 100.0 100.0 100.0
GS 100.0 98.3 100.0 100.0
BG 100.0 100.0 100.0 100.0
Table 10: Accuracy on attack-free samples across different graph topologies on CSQA.ND = No Defense, GS = G-Safeguard, BG = BlindGuard
E.5 Adaptive Attacks
In real-world scenarios, adversaries may learn the principles of a defense mechanism and deliberately adapt their attack strategy to bypass detection. This type of targeted modification is known as an adaptive attack. To evaluate the robustness of BlindGuard under such conditions, we design a dedicated experiment against a simple adaptive Prompt Injection (PI) strategy with GPT-4o-mini as backbone LLM on the PI (CSQA). We modify the original PI attack prompt by embedding the correct answer directly into the input and instructing the agent to “mimic the reasoning process” of the correct answer while explaining the attacker-provided answer, aiming to reduce semantic deviation between malicious responses and normal agent outputs. This strategy is intended to bypass BlindGuard’s anomaly detection. As shown in Table 8, BlindGuard consistently achieved low ASR and the detection AUC remained high, indicating that the hierarchical encoder and corruption-guided detector in BlindGuard were able to capture subtle abnormal interaction patterns and isolate compromised agents, even when their semantic embeddings were close to those of normal agents.
E.6 Encoder Contribution Analysis
To isolate the contribution of our learned encoder, we introduce an additional baseline, Raw SentenceBERT + Eq.(7), which removes the encoder entirely and computes the anomaly score in Eq.(7) directly on non-trained SentenceBERT embeddings. We evaluate this baseline under the same experimental setup (GPT-4o-mini) on CSQA (PI) and CSQA (MA) across four communication topologies (Chain / Random / Star / Tree). As shown in Table 9, directly applying Eq.(7) to raw SBERT features yields consistently inferior performance compared to BlindGuard. The consistent performance gains demonstrate that our encoder training, incorporating semantics-aware directional negatives and margin separation, meaningfully improves the representation geometry, enabling more reliable unsupervised anomaly scoring.
E.7 Attack-free Evaluation
To evaluate whether edge pruning may falsely disrupt normal multi-agent system (MAS) functionality in the absence of attacks, we conduct a no-attack evaluation using the same CSQA (PI) and CSQA (MA) settings with GPT-4o-mini across four communication topologies. As shown in Table 10, BlindGuard does not introduce noticeable utility degradation on clean samples, the accuracy remains essentially unchanged compared to the no-defense baseline, and is comparable to the supervised baseline G-Safeguard. These results indicate that BlindGuard can be safely deployed without significantly disrupting normal MAS functionality.
E.8 Additional Results of BlindGuard
We present more detailed results on CSQA, MMLU, GSM8K, InjecAgent and PoisonRAG benchmarks. As shown in Table 11 and Table 12, additional experiments provide comprehensive multi-round ASR@3 comparisons (R1-R3) across multiple attack types and network topologies, demonstrating BlindGuard’s superior performance over Dominant, PREM, and TAM while approaching G-Safeguard’s effectiveness throughout progressive dialogue stages in various attack types.
To demonstrate BlindGuard’s effectiveness under constrained detection resources where computational resources or operational constraints limit the number of agents that can be simultaneously isolated. The $K=2$ setting represents a cost-effective defense approach, where the system achieves substantial security improvement while minimizing disruption to normal agent operations. As shown in Table 13, BlindGuard shows consistent defense capability, approaching the supervised G-Safeguard’s effectiveness.
PI (CSQA) PI (MMLU) PI (GSM8K)
@@ E.8 Additional Results of BlindGuard
Chain ND 0.31 0.397 0.423 Chain ND 0.283 0.347 0.383 Chain ND 0.064 0.092 0.098
Dom 0.247 0.283 0.303 Dom 0.217 0.240 0.247 Dom 0.071 0.088 0.085
PREM 0.237 0.290 0.293 PREM 0.240 0.257 0.250 PREM 0.061 0.075 0.088
TAM 0.217 0.260 0.273 TAM 0.220 0.233 0.237 TAM 0.064 0.088 0.088
BG 0.203 0.223 0.237 BG 0.190 0.193 0.193 BG 0.061 0.085 0.085
GS 0.187 0.197 0.193 GS 0.190 0.193 0.193 GS 0.044 0.044 0.044
Tree ND 0.293 0.337 0.330 Tree ND 0.250 0.300 0.320 Tree ND 0.075 0.082 0.102
Dom 0.233 0.260 0.270 Dom 0.200 0.200 0.193 Dom 0.061 0.061 0.068
PREM 0.220 0.227 0.267 PREM 0.207 0.220 0.240 PREM 0.064 0.075 0.085
TAM 0.223 0.253 0.260 TAM 0.230 0.233 0.220 TAM 0.071 0.075 0.081
BG 0.227 0.257 0.240 BG 0.207 0.220 0.207 BG 0.075 0.075 0.085
GS 0.180 0.190 0.187 GS 0.187 0.187 0.183 GS 0.071 0.075 0.078
Star ND 0.333 0.453 0.503 Star ND 0.293 0.350 0.390 Star ND 0.071 0.115 0.129
Dom 0.303 0.357 0.387 Dom 0.227 0.257 0.267 Dom 0.054 0.071 0.078
PREM 0.280 0.330 0.343 PREM 0.260 0.297 0.310 PREM 0.068 0.095 0.105
TAM 0.240 0.270 0.287 TAM 0.243 0.263 0.283 TAM 0.061 0.081 0.081
BG 0.193 0.213 0.220 BG 0.200 0.207 0.203 BG 0.068 0.068 0.068
GS 0.183 0.190 0.193 GS 0.177 0.170 0.173 GS 0.054 0.064 0.061
Random ND 0.32 0.377 0.420 Random ND 0.347 0.393 0.410 Random ND 0.059 0.097 0.145
Dom 0.313 0.343 0.370 Dom 0.247 0.280 0.290 Dom 0.054 0.085 0.105
PREM 0.280 0.330 0.347 PREM 0.273 0.337 0.370 PREM 0.037 0.078 0.105
TAM 0.283 0.323 0.353 TAM 0.307 0.380 0.383 TAM 0.047 0.098 0.149
BG 0.243 0.280 0.300 BG 0.220 0.253 0.267 BG 0.041 0.054 0.064
GS 0.190 0.193 0.190 GS 0.180 0.187 0.190 GS 0.048 0.038 0.038
Table 11: Attack success rate (ASR) comparison of defense methods with GPT-4o-mini as backbone LLMs (Part 1). ND = No Defense, Dom = Dominant, BG = BlindGuard, GS = G-Safeguard. Lower values indicate better defense performance.
Appendix F Topology-dependent Performance
We provide an in-depth analysis across topologies showing how topologies mediate attack propagation and BlindGuard’s effectiveness, grounded in the table 1, followed by brief case studies.
Chain.
Propagation is sequential and relatively slow. BlindGuard consistently lowers ASR@3 versus no defense across attacks. The neighbor and global views are complementary in this setting, and local shift is easier to detect, so pruning upstream edges quickly reduces downstream exposure.
Tree.
Subtrees can collude and create strong local consistency while deviating from the global intent. BlindGuard’s system-level context is particularly helpful and delivers large ASR reductions. Across tasks, BlindGuard remains competitive, and AUC indicates reliable separation despite heterogeneous branches.
Star.
An attacked agent can broadcast widely, which makes the topology high risk. BlindGuard benefits from a strong global signal and promptly prunes edges of central agent, yielding marked ASR drops. The approach is effective because central agent misalignment with the global context is easy to detect, and containment at the central agent immediately protects most spokes.
Random.
Low-degree paths dilute neighbor evidence, which makes PI tasks relatively harder to contain, and ASR@3 remains higher than in star or tree. Random connectivity reduces the reliability of neighbor information. The global context stabilizes detection. BlindGuard still achieves strong containment on tool and memory attacks where semantic shift is more uniform.
The star topology has a single center agent that connects to all other agents, and the peripheral agents do not connect to each other. Under most cases of detection errors, an undetected attacked peripheral agent can easily push misleading content to the center, which then distributes it to the entire system. This vulnerability pathway through the center agent is structurally distinct from other topologies, where spread relies on sequential or irregular multi‑path transmission rather than a single global distributor.
[…]
TA (InjecAgent) MA (CSQA) MA (PoisonRAG)
@@ Appendix F Topology-dependent Performance
Chain ND 0.337 0.442 0.480 Chain ND 0.137 0.237 0.267 Chain ND 0.120 0.193 0.223
Dom 0.153 0.142 0.150 Dom 0.137 0.180 0.187 Dom 0.090 0.100 0.110
PREM 0.147 0.152 0.152 PREM 0.083 0.133 0.157 PREM 0.030 0.043 0.060
TAM 0.243 0.293 0.300 TAM 0.090 0.130 0.133 TAM 0.077 0.120 0.147
BG 0.142 0.178 0.164 BG 0.043 0.053 0.070 BG 0.047 0.083 0.100
GS 0.122 0.108 0.102 GS 0.060 0.067 0.077 GS 0.030 0.030 0.040
Tree ND 0.289 0.428 0.451 Tree ND 0.153 0.200 0.213 Tree ND 0.107 0.157 0.180
Dom 0.146 0.149 0.153 Dom 0.150 0.200 0.217 Dom 0.083 0.123 0.143
PREM 0.159 0.182 0.162 PREM 0.087 0.110 0.140 PREM 0.037 0.057 0.083
TAM 0.231 0.289 0.321 TAM 0.113 0.153 0.130 TAM 0.057 0.090 0.120
BG 0.142 0.143 0.125 BG 0.050 0.067 0.077 BG 0.037 0.057 0.070
GS 0.076 0.061 0.048 GS 0.067 0.060 0.067 GS 0.017 0.033 0.030
Star ND 0.368 0.482 0.436 Star ND 0.137 0.207 0.270 Star ND 0.127 0.223 0.260
Dom 0.130 0.151 0.143 Dom 0.133 0.227 0.267 Dom 0.110 0.140 0.137
PREM 0.132 0.139 0.147 PREM 0.100 0.173 0.200 PREM 0.063 0.073 0.090
TAM 0.246 0.266 0.266 TAM 0.110 0.170 0.190 TAM 0.083 0.157 0.197
BG 0.118 0.132 0.126 BG 0.040 0.067 0.093 BG 0.037 0.080 0.100
GS 0.085 0.072 0.069 GS 0.033 0.010 0.017 GS 0.000 0.007 0.017
Random ND 0.336 0.409 0.398 Random ND 0.120 0.213 0.240 Random ND 0.123 0.197 0.247
Dom 0.144 0.160 0.149 Dom 0.147 0.277 0.343 Dom 0.080 0.133 0.170
PREM 0.115 0.136 0.145 PREM 0.100 0.153 0.167 PREM 0.067 0.107 0.137
TAM 0.321 0.383 0.358 TAM 0.150 0.230 0.267 TAM 0.110 0.187 0.237
BG 0.125 0.173 0.177 BG 0.040 0.063 0.087 BG 0.033 0.057 0.067
GS 0.079 0.066 0.051 GS 0.037 0.050 0.050 GS 0.020 0.057 0.070
@@
Table 12: Attack success rate (ASR) comparison of defense methods with GPT-4o-mini as backbone LLMs (Part 2). Abbreviations same as in Table 11.
Appendix G Key Differences from Prior Work
Existing graph anomaly detection (GAD) methods for MAS mainly operate on either structural deviations or static attribute outliers, and are typically designed for fixed graph settings. These methods struggle to detect coordinated semantic anomalies in dynamic, language‑based MAS interactions. Traditional MAS security mechanisms such as rule‑based detection systems rely on rigid, pre‑specified communication formats and lack the capability to handle free‑form, high‑dimensional natural‑language exchanges in MAS. In contrast, BlindGuard introduces a hierarchical agent encoder that jointly captures agent‑level semantics, neighborhood context, and global system state, coupled with a corruption‑guided semantic anomaly detector that learns entirely from benign interactions without any attack labels or prior signatures. This enables BlindGuard to proactively identify and isolate malicious agents within the message‑passing process, blocking adversarial propagation before it contaminates other agents, and distinguishing it from purely structural detection, output‑level consensus, or rule‑driven MAS security designs.
Appendix H Directional Negative Justification
A potential concern is whether BlindGuard relies on isotropic corruption as its primary modeling assumption. We clarify that this is not the case. Our defense is built around semantics-aware directional negatives and an angular-margin separation from the normal manifold, which is more closely aligned with how real semantic attacks deviate from normal behavior. We encode each agent’s message into an embedding and compute global (system-level) and neighborhood context vectors. The anomaly score measures alignment with these contexts via cosine similarity. Real semantic attacks tend to shift an agent’s message away from the round’s intent, reducing alignment with global and neighborhood contexts. Training with directional negatives enforces an angular (cosine) margin around the normal manifold. Any perturbation whose direction falls within a cone opposing the context vectors must reduce alignment beyond this margin, and thus becomes detectable. This mechanism is standard in metric and contrastive learning: semantics-aware hard negatives tighten the decision boundary precisely in the regions where real attacks are most likely to occur, whereas isotropic random noise does not provide such targeted boundary refinement. We acknowledge, however, that in stealthier attack scenarios where adversarial perturbations are carefully crafted to remain close to the normal manifold, the angular margin may be insufficient, and we leave more robust defenses against such scenarios for future work.
Random Topology Chain Topology
Method R1 R2 R3 Method R1 R2 R3
No Defense 0.3467 0.3933 0.4100 No Defense 0.2833 0.3467 0.3833
Dominant 0.2933 0.3767 0.4000 Dominant 0.2233 0.2367 0.2600
PREM 0.3100 0.3933 0.4300 PREM 0.2467 0.2667 0.2767
TAM 0.3267 0.3833 0.3933 TAM 0.2267 0.2600 0.2733
BlindGuard 0.2433 0.2967 0.3400 BlindGuard 0.2467 0.2667 0.2667
G-Safeguard 0.2200 0.2500 0.2633 G-Safeguard 0.2133 0.2400 0.2467
Star Topology Tree Topology
Method R1 R2 R3 Method R1 R2 R3
No Defense 0.2933 0.3500 0.3900 No Defense 0.2500 0.3000 0.3200
Dominant 0.2700 0.3200 0.3500 Dominant 0.2333 0.2467 0.2600
PREM 0.2667 0.3133 0.3367 PREM 0.2267 0.2567 0.2700
TAM 0.2467 0.2700 0.2767 TAM 0.2067 0.2233 0.2200
BlindGuard 0.2233 0.2467 0.2733 BlindGuard 0.2100 0.2267 0.2267
G-Safeguard 0.2300 0.2567 0.2700 G-Safeguard 0.2100 0.2200 0.2200
Table 13: Attack success rate (ASR) across communication rounds with detection budget $K=2$.
Appendix I Failure Case Analysis
We examine a representative query from our dataset: “Marcus is trying to decide whether he really needs to do his homework… What is the percentage chance that Marcus will actually have to turn in his homework tomorrow?” From this case, we find that agents can legitimately adopt different reasoning styles, and this style diversity helps explain both types of BlindGuard errors we observed. Normal agents use atypical direct-computation formats or introduce extra heuristics, creating local/global semantic disagreement and being flagged as outliers.
Why a “real-world nuance” style can cause false negatives (malicious not flagged)? Some malicious agents answer by staying perfectly on-topic while adding plausible-sounding caveats, e.g., “substitutes often still collect homework,” “teacher mood matters,” “events aren’t independent.” This is not an off-topic derailment; it is manifold-preserving mimicry. Semantically, the message heavily overlaps with the round intent (homework, substitute, extension, probabilities), so its embedding can remain highly aligned with the global/neighborhood context. Since BlindGuard’s anomaly score mainly relies on semantic misalignment to context, such on-topic, stylistically consistent steering may not reduce alignment enough to cross the detection margin, leading to a miss.
Why a “direct computation” style can cause false positives (benign flagged)? Other benign agents follow the expected probabilistic calculation strictly and produce a concise numeric derivation. In contrast, some benign replies may introduce extra heuristics (e.g., “assume the substitute collects homework with 50% chance”) or use a markedly different structured format. Even if non-malicious, this can create local/global semantic disagreement with the dominant group reasoning pattern for that round, making the agent look like an outlier under a context-alignment detector and thus potentially triggering false pruning.
These cases show BlindGuard is strongest at detecting semantic deviation (off-manifold behavior), but may struggle with (1) stealthy, on-topic critique attacks and (2) benign heterogeneity in reasoning style.